# Multivariate Fields of Experts

### v1

@@ IV-C2 Compressed-Sensing MRI
Zero-fill 27.43 29.62 23.49 27.00
TV 33.23 32.70 26.97 29.42
WCRR 35.10 34.18 29.32 31.00
MFoE 35.40 34.21 30.26 31.20
Prox-DRUNet 35.32 34.53 30.97 31.66
@@
IV-C3 Computed Tomography
We evaluate our method on the LoDoPaB-CT dataset [33], which contains ground-truth images derived from 2D slices of reconstructions in the LIDC/IDRI database [34], each of size $\left(362\times 362\right)$. The measurement operators are defined by parallel-beam geometry and implemented using the ASTRA Toolbox [35, 36]. All experiments are conducted with 256 simulated detectors. The resulting sinograms are corrupted with Gaussian noise of standard deviation $\sigma_{\mathbf{w}}=0.1$.
We consider three acquisition settings with 60, 40, and 20 projection angles, respectively. Each method is initialized with the filtered-back projected (FBP) image, and the hyperparameters are tuned on a validation set of 10 images. The performance is reported on a testing set of 50 images, all normalized to the range $[0,1]$.
@@ IV-C3 Computed Tomography
TABLE V: PSNR values for CT reconstruction on the LoDoBaP dataset.
@@
FBP 28.53 25.87 21.41
TV 33.74 32.58 30.11
WCRR 35.21 33.87 31.08
MFoE 35.45 34.22 31.54
Prox-DRUNet 35.73 34.61 32.12
@@
IV-C4 Time Comparisons
Table VI contains the average inference time per image, measured in seconds, for each inverse problem. More detailed results are provided in Tables $X$. These results highlight that the performance gains of Prox-DRUNet come at a computational cost. In particular, the CT reconstructions with Prox-DRUNet are significantly slower than with other methods, as they require the repeated evaluation of the proximal operator of the data-fitting term—a step that is especially costly in the case of CT. By contrast, MFoE is only moderately more demanding than the WCRR in inference time, between 35% and 40% on average across all experiments due to the computational cost of the projection onto $\mathcal{B}_{\ell_{1}}$.
TABLE VI: Average time per image for all the reconstruction modalities.
@@ IV-C4 Time Comparisons
Deep-learning-based methods such as Prox-DRUNet sets a high standard in learned regularization by leveraging a large number of parameters and extensive training data which contribute to their good performance across diverse tasks. We believe that MFoE offers an interesting alternative with significantly fewer parameters, fewer training data, faster inference time, and with a reconstruction quality that is close to Prox-DRUNet. Together, these methods offer complementary strengths, providing valuable options suited to varying application needs.
V Conclusion
We have proposed the MFoE model. It is a principled extension of the classic FoE model and incorporates multivariate potential functions via Moreau envelopes of the $\ell_{\infty}$-norm. Our method offers improved adaptability across multiple inverse problems. The image quality that it achieves is comparable to that of state-of-the-art learned regularizers. It offers a clear structured design that is easy to implement with competitive reconstruction times.
@@ Acknowledgements
The research leading to these results was supported by the European Research Council (ERC) under European Union’s Horizon 2020 (H2020), Grant Agreement - Project No 101020573 FunLearn.

### v2

@@ VI-B Compressed-Sensing MRI
Zero-fill 27.43 / 0.843 29.62 / 0.875 23.49 / 0.764 27.00 / 0.833
TV 33.23 / 0.920 32.70 / 0.905 26.97 / 0.835 29.42 / 0.858
WCRR 35.10 / 0.950 34.18 / 0.928 29.32 / 0.892 31.00 / 0.888
MFoE 35.40 / 0.952 34.21 / 0.928 30.22 / 0.903 31.21 / 0.888
Prox-DRUNet 35.32 / 0.950 34.53 / 0.931 30.97 / 0.904 31.66 / 0.894
@@
VI-C Computed Tomography
We evaluate our method on the LoDoPaB-CT dataset [23], which contains ground-truth images derived from 2D slices of reconstructions in the LIDC/IDRI database [3], each one of size $\left(362\times 362\right)$. The measurement operators are defined by parallel-beam geometry and implemented using the ASTRA Toolbox [42, 29]. All experiments are conducted with 256 simulated detectors. The resulting sinograms are corrupted with Gaussian noise of standard deviation $\sigma_{\mathbf{w}}=0.1$.
We consider three acquisition settings with 60, 40, and 20 projection angles, respectively. Each method is initialized with the filtered-back-projected (FBP) image, and the hyperparameters are tuned on a validation set of 10 images. The performance is reported on a testing set of 50 images, all normalized to the range $[0,1]$.
@@ VI-C Computed Tomography
TABLE V: Quantitative results (PSNR / SSIM) for CT reconstruction on the LoDoBaP dataset.
@@
FBP 28.53 / 0.796 25.87 / 0.691 21.41 / 0.494
TV 33.74 / 0.917 32.58 / 0.903 30.11 / 0.867
WCRR 35.21 / 0.924 33.87 / 0.912 31.08 / 0.880
MFoE 35.40 / 0.925 34.21 / 0.914 31.48 / 0.885
Prox-DRUNet 35.73 / 0.927 34.61 / 0.917 32.12 / 0.893
@@
VI-D Statistical Significance
Across all our inverse problem experiments, we observe that WCRR outperforms TV, MFoE outperforms WCRR, and Prox-DRUNet usually outperforms MFoE. We assess the validity of this ranking through Wilcoxon signed-rank tests [44] on the PSNR and SSIM scores across the entire test set, with a strict significance threshold of $p<0.001$. In total, we perform 26 separate tests, covering the PSNR and SSIM metrics across all 13 experimental modalities (six deblurring, four CS-MRI, and three CT).
Our testing framework confirms the robustness of this hierarchy. WCRR yields statistically significant improvements over TV in every single configuration. The transition from WCRR to MFoE is similarly distinct, with MFoE surpassing WCRR in all PSNR evaluations and achieving significant SSIM gains in 10 out of 13 modalities. Finally, Prox-DRUNet proves superior to MFoE in the majority of scenarios; statistical significance is established in 19 of the 26 comparisons, with the few exceptions occurring primarily in deblurring and CS-MRI tasks where both models achieved comparable high-fidelity reconstructions.
VI-E Time Comparisons
Table VI contains the average inference time per image, measured in seconds, for each inverse problem. More detailed results are provided in Tables $X$. These results highlight that the performance gains of Prox-DRUNet come at a computational cost. In particular, the CT reconstructions with Prox-DRUNet require the repeated evaluation of the proximal operator of the data-fitting term—a step that is especially costly in the case of CT. This results in Prox-DRUNet being significantly slower than the other methods. The MFoE and WCRR methods are on average more than 13 times faster than Prox-DRUNet.
TABLE VI: Average duration [s] per image for all the reconstruction modalities.
@@ VI-E Time Comparisons
Although deep-learning-based methods such as Prox-DRUNet set a high standard in learned regularization, they do so at the cost of a large number of parameters and extensive training data. We believe that MFoE offers an interesting alternative with significantly fewer parameters, fewer training data, faster inference time, and with a reconstruction quality that is close to Prox-DRUNet. Together, these methods offer complementary strength and provide valuable options suited to a variety of application needs.
VII Conclusion
We have proposed the MFoE model. It is a principled extension of the classic FoE model and incorporates multivariate potential functions via Moreau envelopes of the $\ell_{\infty}$-norm. The image quality that it achieves noticeably narrows the performance gap with state-of-the-art learned regularizers while it offers a clear structured design that is easy to implement with competitive reconstruction times.
@@ Acknowledgements
The research leading to these results was supported by the European Research Council (ERC) under European Union’s Horizon 2020 (H2020), Grant Agreement - Project No. 101020573 FunLearn.